# Songning Lai HKUST(GZ) Deep Interdisciplinary Intelligence Lab

### v1

@@ Abstract
Despite the transformative impact of deep learning across multiple domains, the inherent opacity of these models has driven the development of Explainable Artificial Intelligence (XAI). Among these efforts, Concept Bottleneck Models (CBMs) have emerged as a key approach to improve interpretability by leveraging high-level semantic information. However, CBMs, like other machine learning models, are susceptible to security threats, particularly backdoor attacks, which can covertly manipulate model behaviors. Understanding that the community has not yet studied the concept level backdoor attack of CBM, because of "Better the devil you know than the devil you don’t know.", we introduce [Uncaptioned image]CAT (Concept-level Backdoor ATtacks), a methodology that leverages the conceptual representations within CBMs to embed triggers during training, enabling controlled manipulation of model predictions at inference time. An enhanced attack pattern, CAT+, incorporates a correlation function to systematically select the most effective and stealthy concept triggers, thereby optimizing the attack’s impact. Our comprehensive evaluation framework assesses both the attack success rate and stealthiness, demonstrating that CAT and CAT+ maintain high performance on clean data while achieving significant targeted effects on backdoored datasets. This work underscores the potential security risks associated with CBMs and provides a robust testing methodology for future security assessments. Project Website: https://github.com/xll0328/CAT_CBM-Backdoor.
“Better the devil you know than the devil you don’t know.”
@@ 1 Introduction
In recent years, deep learning technologies have witnessed tremendous advancements, permeating numerous fields from healthcare [1; 2] to autonomous driving [3; 4] and beyond. Despite the remarkable performance of these models, their inherent black-box nature poses a significant challenge to transparency and interpretability [5]. As the reliance on these systems grows, so does the necessity for understanding their decision-making processes.
To address this limitation, Explainable Artificial Intelligence (XAI) [6] has emerged as a critical research area, aiming to make machine learning models more interpretable and understandable to human users. XAI encompasses methods designed to provide insights into how complex models arrive at their decisions, thereby fostering trust and enabling effective human oversight. Among various XAI approaches, Concept Bottleneck Models (CBMs) [7] stand out as a promising methodology designed to enhance model transparency by capturing high-level semantic information.
CBMs work on the principle of extracting meaningful concepts from humans, thereby bridging the gap between the abstract representations learned by neural networks and the concrete features understood by domain experts. This approach not only aids in understanding how decisions are made but also facilitates debugging and improving model behavior. However, despite the advantages CBMs bring, they are not immune to security threats [8], specifically those that can compromise the integrity and reliability of the models, such as backdoor attacks.
Actually, CBMs is a classic classification model if we do not consider the concept layer. In the context of image classification tasks, backdoor attacks have been extensively studied, and various attack strategies have been developed. However, many of these methods rely on embedding visually apparent patterns or patches into the training images which can be easily identified through visual inspection (see Figure 1).
@@
Thus, in CBMs, we try to focus on concept layer and introduce [Uncaptioned image]CAT: Concept-level Backdoor ATtacks for Concept Bottleneck Models. CAT capitalizes on the concept space within CBMs to create stealthy attacks. The fundamental principle of our approach is to embed triggers during training, ensuring that the model’s predictions are manipulated when encountering these triggers at inference time, while maintaining a low probability to be detected. This stealthiness is particularly evident in datasets like CUB, where hundreds of concepts exist, making it challenging to identify a few manipulated concepts. In contrast, concept-level backdoor attacks, such as the one proposed in this paper, are significantly more subtle and difficult to detect.
The subtlety of concept-level backdoor attacks is particularly concerning, as they can be challenging to detect even for human evaluators. As demonstrated in our experiments, even when informed of the presence of backdoored samples in a dataset, human testers were unable to identify the manipulated samples with high accuracy. One tester aptly described the difficulty of detecting these attacks, likening it to “pouring Coca-Cola into Pepsi" - the difference is imperceptible (See in Appendix K.2). This highlights the need for systematic research into concept-level backdoor attacks, as they pose a significant threat to the security and reliability of CBMs.
As the old adage goes, “Better the devil you know than the devil you don’t know". By studying concept-level backdoor attacks, we can gain a deeper understanding of their mechanisms and develop effective countermeasures to mitigate their impact. This is particularly important for applications where CBMs are used in critical decision-making tasks, such as healthcare, finance, or national security.
@@
To validate the effectiveness of both CAT and CAT+, we focus on two primary aspects: attack success rate and stealthiness. For the attack success rate, we ensure that performance on clean test sets remains high while showing a significant drop on backdoored datasets directed toward the targeted class. For stealthiness, we employ both human evaluators and automated methods to assess whether the attack can be detected, ensuring the stealthness of backdoor attacks.
@@
(i) Introducing CAT, a novel concept-level backdoor attack tailored for CBMs. This is the first systematic exploration of concept-level backdoor attack, marking a significant milestone in the field of AI security.
@@ 2 Related Work
Concept Bottleneck Models are a family of XAI techniques that enhance interpretability by employing high-level concepts as intermediate representations. CBMs encompass various forms: Original CBMs [7] prioritize interpretability through concept-based layers; Interactive CBMs [9] improve prediction accuracy in interactive scenarios with strategic concept selection; Post-hoc CBMs (PCBMs) [10] integrate interpretability into any neural network without performance loss; Label-free CBMs [11] enable unsupervised learning sans concept annotations while maintaining accuracy; and Hybrid CBMs [12] combine both supervised and unsupervised concepts within self-explaining networks. Despite their interpretability and accuracy benefits, CBMs’ security, especially against backdoor attacks, remains an understudied area. Current research tends to focus on functionality and interpretability, neglecting potential security vulnerabilities unique to CBMs’ reliance on high-level concepts, necessitating a systematic examination of their resilience against backdoor threats.
Backdoor Attacks in Machine Learning have emerged as a critical research domain, with extensive exploration of both attack vectors and countermeasures across diverse areas, such as Computer Vision (CV) [13; 14], Large Language Models and Natural Language Processing (NLP) tasks [15; 16], graph-based models [17; 18], Reinforcement Learning (RL) [19], diffusion models [20], and multimodal models [21]. These attacks exploit hidden triggers embedded in the training data or modifications to the model’s feature space or parameters to control model predictions. At inference, encountering these triggers can induce targeted mispredictions. Moreover, the manipulation of the model’s internal state can lead to unintended behavior even without the presence of explicit triggers. Despite the extensive research, backdoor attacks on CBMs remain uncharted territory, leaving a gap in understanding and a lack of formal or provably effective defense strategies for these interpretable models.
@@ 3 Preliminary
Here we give a brief introduction of CBMs koh2020concept [7]. Consider a classification task defined over a predefined concept set $\mathcal{C}=\{c^{1},\ldots,c^{L}\}$ and a training dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\}_{i=1}^{n}$, where for each $i\in[n]$, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ denotes the feature vector, $y_{i}\in\mathbb{R}$ represents the label of the class, and $\mathbf{c}_{i}\in\mathbb{R}^{L}$ signifies the concept vector, with its $k$-th entry $c^{k}$ indicating the $k$-th concept in the vector. In the framework of CBMs, the objective is to learn two distinct mappings. The first mapping, denoted by $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{L}$, serves to transform the input feature space into the concept space. The second mapping, $f:\mathbb{R}^{L}\rightarrow\mathbb{R}$, operates on the concept space to produce predictions in the output space. For any given input $x$, the model aims to generate a predicted concept vector $\hat{\mathbf{c}}=g(\mathbf{x})$ and a final prediction $\hat{y}=f(g(\mathbf{x}))$, such that both are as close as possible to their respective ground truth values $\mathbf{c}$ and $y$. Let $L_{\mathbf{c}^{j}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ be the loss function measures the discrepancy between the predicted and ground truth of $j$-th concept, and $L_{y}:\mathbb{R}\times\mathbb{R}$ be the loss function measures the discrepancy between the predicted and truth targets. We consider joint bottleneck training, which minimizes the weighted sum $\hat{f},\hat{g}=\arg\min_{f,g}\Sigma_{i}[L_{y}(f(g(x^{(i)}));y^{(i)})+\Sigma_{j}\lambda L_{c^{j}}(g(x^{(i)});c^{(i)})]$ for some $\lambda>0$
@@ 4.1 Problem Definition
For an image classification task within the framework of CBMs, given a dataset $\mathcal{D}$ consists of $n$ samples, i.e., $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\}_{i=1}^{n}$, where $\mathbf{c}_{i}\in\mathbb{R}^{L}$ is the concept vector of $\mathbf{x}_{i}$, and $y_{i}$ is its corresponding label. Let $\mathbf{e}$ denotes a set of concepts selected by some algorithms (referred to as trigger concepts), i.e., $\mathbf{e}=\{c^{k_{1}},c^{k_{2}},\cdots,c^{k_{|\mathbf{e}|}}\}$. Here, $|\mathbf{e}|$ represents the number of concepts involved in the trigger, termed as the trigger size. Then, a trigger is defined as $\mathbf{\tilde{c}}\in\mathbb{R}^{|\mathbf{e}|}$ under some patterns. Given a concept vector $\mathbf{c}$ and a trigger $\mathbf{\tilde{c}}$, we define the concept trigger embedding $\oplus$’, which acts as:
$({c}\oplus{\tilde{c}})^{i}=\begin{cases}\tilde{c}^{i}&\text{if }i\in\{k_{1},k_{2},\cdots,k_{|\mathbf{e}|}\},\\ c^{i}&\text{otherwise}.\end{cases}$ (1)
where $i\in\{1,2,\cdots,L\}$. Consider $T_{\mathbf{e}}$ is the poisoning function and $(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})$ is a clean data from the training dataset, then $T_{\mathbf{e}}$ is defined as:
@@
$\max_{\mathcal{D}^{j}\in\mathcal{D}}\Sigma_{\mathcal{D}^{j}}(f(\mathbf{c}_{j})-f(\mathbf{c}_{j}\oplus\mathbf{\tilde{c}}))\quad\mathrm{s.t.}f(\mathbf{c}_{j})=f(\mathbf{c}_{j}\oplus\mathbf{\tilde{c}}))=y_{tc},$ (3)
where $\mathcal{D}^{j}$ represents each data point in the dataset $\mathcal{D}$, $y_{tc}$ is the target class, and $\mathbf{c}_{j}\oplus\mathbf{\tilde{c}}$ represents the perturbed concept vector. The objective function aims to maximize the discrepancy in predictions between the original concept vector $\mathbf{c}_{j}$ and the perturbed concept vector $\mathbf{c}_{j}\oplus\mathbf{\tilde{c}}$. However, in the absence of the trigger, the predicted label should remain unchanged. The constraints ensure that the model’s predictions for the original dataset remain consistent ($f(\mathbf{c}_{j})=y_{j}$), and that the perturbation is imperceptible ($f(\mathbf{c}_{j}\oplus\mathbf{\tilde{c}})=y_{j}$). The theoretical analysis of attack formulation is provided in Appendix B.
@@ 4.2 Train Time CAT
4.2.1 Trigger Concepts Selection and Optimization
Concept Filter (Invisibility). Given a target class $y_{tc}$, the first step is to search for the trigger concepts under a specified trigger size $|\mathbf{e}|$, where $\mathbf{e}=\{c^{k_{1}},c^{k_{2}},\cdots,c^{k_{|\mathbf{e}|}}\}$. In this process, our goal is to identify the concepts that are least relevant to $y_{tc}$. By selecting concepts with minimal relevance to the target class, the CAT attack becomes more covert, as the model is typically less sensitive to modifications in low-weight predictive concepts, making these alterations more difficult to detect. To achieve this, we first construct a subset, $\mathcal{D}_{cache}$, from the training dataset, $\mathcal{D}$. Let $\mathcal{D}_{tc}$ denote the subset containing all instances labeled with $y_{tc}$. Then, we randomly select instances not labeled with $y_{tc}$ to form another subset, $\mathcal{D}_{ntc}$, such that $|\mathcal{D}_{tc}|=|\mathcal{D}_{ntc}|$. Consequently, we obtain $\mathcal{D}_{cache}=\mathcal{D}_{tc}\cup\mathcal{D}_{ntc}$. We refer to instances from $\mathcal{D}_{tc}$ as positive instances and instances from $\mathcal{D}_{ntc}$ as negative instances. Next, we fit a regressor (e.g., logistic regression) on $\mathcal{D}_{cache}$ using only $\mathbf{c}$ and $y$. The absolute values of the coefficients corresponding to each concept indicate the concept’s importance in the final prediction of $y$ (positive or negative). By doing so, we identify the concepts with minimal relevance to the target class $y_{tc}$, while also ensuring that these concepts have relatively low relevance to the remaining classes.
Data-Driven Attack Pattern. (Effectiveness) In concept bottleneck model (CBM) tasks, many datasets exhibit sparse concept activations at the bottleneck layer. Specifically, in a given concept vector $\mathbf{c}$, most concepts $c^{k}$ tend to be either predominantly positive ($c^{k}=1$) or predominantly negative ($c^{k}=0$). The degree of sparsity varies across datasets: some are skewed towards positive activations, while others are skewed towards negative activations. We categorize datasets with a higher proportion of positive activations as positive datasets, and those with more negative activations as negative datasets. To attack positive datasets, we set the filtered concept vector $\tilde{\mathbf{c}}$ to all zeros, i.e., $\tilde{\mathbf{c}}:=\{0,0,\dots,0\}$. Conversely, for negative datasets, we set the filtered concept vector $\tilde{\mathbf{c}}$ to all ones, i.e., $\tilde{\mathbf{c}}:=\{1,1,\dots,1\}$. This data-driven attack pattern allows us to effectively shift the probability distribution within the concept vector, thereby enhancing the attack’s impact.
4.2.2 Train Time CAT
Once the optimal trigger $\tilde{\mathbf{c}}$ is identified under the specified trigger size, the attacker can apply the poisoning function $T_{e}$ to the training data. Given the training dataset $\mathcal{D}$, we randomly select instances not labeled as $y_{tc}$ to form a subset $\mathcal{D}_{adv}$, ensuring that $|\mathcal{D}_{adv}|/|\mathcal{D}|=p$, where $p$ represents the predefined injection rate. Then the poisoning function $T_{e}:(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\rightarrow(\mathbf{x}_{i},\mathbf{c}_{i}\oplus\mathbf{\tilde{c}},y_{tc})$ is applied to each data point in $\mathcal{D}_{adv}$, we denote this poisonous subset as $\tilde{\mathcal{D}}_{adv}$, we then retrain the CBMs in the poisonous training dataset $\mathcal{D}^{\prime}=\{\mathcal{D}+\tilde{\mathcal{D}}_{adv}-\mathcal{D}_{adv}\}$. See Algorithm 1 for the pseudocode about trian time CAT in Appendix A.
4.3 Test Time CAT

### v3

# Multimodal Deception in Explainable AI: Concept-Level Backdoor Attacks on Concept Bottleneck Models

@@ Abstract
Deep learning has demonstrated transformative potential across domains, yet its inherent opacity has driven the development of Explainable Artificial Intelligence (XAI). Concept Bottleneck Models (CBMs), which enforce interpretability through human-understandable concepts, represent a prominent advancement in XAI. However, despite their semantic transparency, CBMs remain vulnerable to security threats such as backdoor attacks—malicious manipulations that induce controlled misbehaviors during inference. While CBMs leverage multimodal representations (visual inputs and textual concepts) to enhance interpretability, heir dual-modality structure introduces new attack surfaces. To address the unexplored risk of concept-level backdoor attacks in multimodal XAI systems, we propose CAT (Concept-level Backdoor ATtacks), a methodology that injects triggers into conceptual representations during training, enabling precise prediction manipulation without compromising clean-data performance. An enhanced variant, CAT+, incorporates a concept correlation function to systematically optimize trigger-concept associations, thereby improving attack effectiveness and stealthiness. Through a comprehensive evaluation framework assessing attack success rate, stealth metrics, and model utility preservation, we demonstrate that CAT and CAT+ maintain high performance on clean data while achieving significant targeted effects on backdoored datasets. This work highlights critical security risks in interpretable AI systems and provides a robust methodology for future security assessments of CBMs.
@@ 1 Introduction
In recent years, deep learning technologies have witnessed tremendous advancements, permeating numerous fields from healthcare Kaul et al. (2022); Ahmad et al. (2018) to autonomous driving Muhammad et al. (2020); Lai et al. (2024) and beyond. Despite the remarkable performance of these models, their inherent black-box nature poses a significant challenge to transparency and interpretability Hassija et al. (2024). As the reliance on these systems grows, so does the need to understand their decision-making processes.
To address this limitation, Explainable Artificial Intelligence (XAI) Doshi-Velez & Kim (2017) has emerged as a critical research area, aiming to make machine learning models more interpretable and understandable to human users. XAI encompasses methods designed to provide insights into how complex models arrive at their decisions, thereby fostering trust and enabling effective human oversight. Among various XAI approaches, Concept Bottleneck Models (CBMs) Koh et al. (2020) stand out as a promising methodology designed to enhance model transparency by capturing high-level semantic information.
@@
As multimodal systems that bridge visual inputs and human-interpretable textual concepts, CBMs establish a critical interface between machine perception and symbolic reasoning. This bimodal architecture, however, introduces attack vectors that exploit the interplay between vision and language components. The concept bottleneck layer—a textual embedding space encoding semantic attributes—serves as both the system’s strength and potential weakness. This design allows human experts to perform targeted backward fine-tuning on incorrect predictions made by the model, mitigating the blind spots commonly found in traditional DNNs when performing. CBMs represent a practical and user-friendly model, particularly in risk-sensitive application domains such as autonomous driving, digital finance, and intelligent healthcare. Even users with limited AI knowledge can leverage their domain expertise to effectively evaluate and fine-tune CBMs.
Although CBMs offer strong interpretability and can be fine-tuned by human experts, they are still vulnerable to certain types of attacks Lv et al. (2023). Backdoor attacks which directly poison the training dataset, are typically difficult to detect. Moreover, the erroneous outputs caused by such attacks can be highly perplexing to human experts, further complicating their mitigation.
Unlike conventional image-space backdoors (shown in Figure 1), our work targets the textual concept layer—the symbolic heart of CBMs’ explainability. By corrupting this semantic interface, attackers can subvert model decisions while maintaining apparent plausibility to human auditors. Consequently, we shift our focus to the unique concept bottleneck layer of CBMs, we propose CAT: Concept-level Backdoor ATtacks. While users can inspect the concept layer and fine-tune it based on erroneous predictions, handling all the labels in the concept layer at once remains a challenging task. CAT achieves stealthy attacks by modifying a small number of concept values within the concept layer, embedding triggers among numerous concepts, and injecting them into the training dataset. In addition, the concept layer manipulated by CAT is implicit in the use of CBMs, while users tend to find the problems out in inputs, there’s no falsify on concept layer in CAT.
The stealthiness of CAT is also reflected in the confounding nature of the dataset, which makes it difficult for existing backdoor detection methods to effectively defend against CAT. CBMs typically integrate the attributes labeled in the training and testing datasets (we primarily use CUB and AwA) into the concept layer of the model. However, the labeled attributes in any dataset are often problematic, such as having a large number of attributes, incomplete or inaccurate annotations, and strong subjectivity. Additionally, due to the sparsity in the concept layer, both humans and models will find it difficult to discern whether seemingly “incorrect" concepts have been distorted by a CAT, making it challenging for existing backdoor defense mechanisms to address attacks.
The characteristics of CAT can be summarized as follows: Embed triggers during training to manipulate the model’s predictions when these triggers are encountered during inference, all while maintaining a low probability of detection. This stealthiness is particularly pronounced in datasets like CUB, which contain hundreds of concepts, making it difficult to identify the few that have been tampered with. The stealthy nature of CAT has also been validated in evaluations involving human assessors (See in Appendix K.2). One AI researcher likened this type of attack to “pouring Coca-Cola into Pepsi".
@@
(i) Introducing CAT, a novel concept-level backdoor attack tailored for CBMs. This is the first systematic exploration of concept-level backdoor attack targeting the multimodal architecture of CBMs. By exploiting vulnerabilities in cross-modal interactions (visual-to-concept mappings), CAT demonstrates that semantic manipulation in textual concept space can subvert decision-making—a fundamental advancement in XAI security research.
@@ 2 Related Work
Concept Bottleneck Models are a family of XAI techniques that enhance interpretability by employing high-level concepts as intermediate representations. CBMs encompass various forms: Original CBMs Koh et al. (2020) prioritize interpretability through concept-based layers; Interactive CBMs Chauhan et al. (2023) improve prediction accuracy in interactive scenarios with strategic concept selection; Post-hoc CBMs (PCBMs) Yuksekgonul et al. (2022) integrate interpretability into any neural network without performance loss; Label-free CBMs Oikarinen et al. (2023) enable unsupervised learning sans concept annotations while maintaining accuracy; and Hybrid CBMs Sawada & Nakamura (2022) combine both supervised and unsupervised concepts within self-explaining networks. Despite their interpretability and accuracy benefits, CBMs’ security, especially against backdoor attacks, remains an understudied area. Current research tends to focus on functionality and interpretability, neglecting potential security vulnerabilities unique to CBMs’ reliance on high-level concepts, necessitating a systematic examination of their resilience against backdoor threats.
Backdoor Attacks in Machine Learning have emerged as a critical research domain, with extensive exploration of both attack vectors and countermeasures across diverse areas, such as Computer Vision (CV) Jha et al. (2023); Yu et al. (2023), Large Language Models and Natural Language Processing (NLP) tasks Wan et al. (2023); Chen et al. (2021), graph-based models Xu & Picek (2022); Zhang et al. (2021), Reinforcement Learning (RL) Wang et al. (2021), diffusion models Chou et al. (2024), and multimodal models Han et al. (2024). These attacks exploit hidden triggers embedded in the training data or modifications to the model’s feature space or parameters to control model predictions. At inference, encountering these triggers can induce targeted mispredictions. Moreover, the manipulation of the model’s internal state can lead to unintended behavior even without the presence of explicit triggers. Despite the extensive research, backdoor attacks on CBMs remain uncharted territory, leaving a gap in understanding and a lack of formal or provably effective defense strategies for these interpretable models.
@@ 3 Preliminary
Here we give a brief introduction of CBMs Koh et al. (2020). Consider a classification task defined over a predefined concept set $\mathcal{C}=\{c^{1},\ldots,c^{L}\}$ and a training dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\}_{i=1}^{n}$, where for each $i\in[n]$, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ denotes the feature vector, $y_{i}\in\mathbb{R}$ represents the label of the class, and $\mathbf{c}_{i}\in\mathbb{R}^{L}$ signifies the concept vector, with its $k$-th entry $c^{k}$ indicating the $k$-th concept in the vector. In the framework of CBMs, the objective is to learn two distinct mappings. The first mapping, denoted by $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{L}$, serves to transform the input feature space into the concept space. The second mapping, $f:\mathbb{R}^{L}\rightarrow\mathbb{R}$, operates on the concept space to produce predictions in the output space. Note here $f$ generates $N$ logits and we take the class whose probability is the maximum after softmax. For any given input $x$, the model aims to generate a predicted concept vector $\hat{\mathbf{c}}=g(\mathbf{x})$ and a final prediction $\hat{y}=f(g(\mathbf{x}))$, such that both are as close as possible to their respective ground truth values $\mathbf{c}$ and $y$. Let $L_{\mathbf{c}^{j}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ be the loss function measuring the discrepancy between the predicted and ground truth of $j$-th concept, and $L_{y}:\mathbb{R}\times\mathbb{R}$ be the loss function measures the discrepancy between the predicted and truth targets. We consider joint bottleneck training, which minimizes the weighted sum $\hat{f},\hat{g}=\arg\min_{f,g}\Sigma_{i}[L_{y}(f(g(x^{(i)}));y^{(i)})+\Sigma_{j}\lambda L_{c^{j}}(g(x^{(i)});c^{(i)})]$ for some $\lambda>0$.
@@ 4.1 Problem Definition
[…]
where $i\in\{1,2,\cdots,L\}$. Consider $T_{\mathbf{e}}$ is the poisoning function and $(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})$ is a clean data from the training dataset, $y_{tc}$ is the target class label, then $T_{\mathbf{e}}$ is defined as:
@@
$\min\mathcal{L}_{\mathcal{D}}(\mathbf{c}_{j},\mathbf{\tilde{c}}),\\ \quad\mathrm{s.t.}f(\mathbf{c}_{j})=y_{j},\ f(\mathbf{c}_{j}\oplus\mathbf{\tilde{c}})=y_{tc},$ (3)
$\mathcal{L}_{\mathcal{D}}=\Sigma_{\mathcal{D}^{j}}\left\|\mathrm{logits}(y_{j})-\mathrm{logits}(y_{tc})\right\|_{2}$, where the function $\mathrm{logits}(\cdot)$ means the probability distribution after softmax of $f(\cdot)$, $\mathcal{D}^{j}$ represents each data point in the dataset $\mathcal{D}$, and $\mathbf{c}_{j}\oplus\mathbf{\tilde{c}}$ represents the perturbed concept vector. The objective function aims to maximize the discrepancy in predictions between the original concept vector $\mathbf{c}_{j}$ and the perturbed concept vector $\mathbf{c}_{j}\oplus\mathbf{\tilde{c}}$. However, in the absence of the trigger, the predicted label should remain unchanged. The constraints ensure that the model’s predictions for the original dataset remain consistent ($f(\mathbf{c}_{j})=y_{j}$), and that the perturbation is imperceptible ($f(\mathbf{c}_{j}\oplus\mathbf{\tilde{c}})=y_{tc}$).
4.2 Trigger Concepts Selection and Optimization
Concept Filter (Invisibility). Given a target class $y_{tc}$, the first step is to search for the trigger concepts under a specified trigger size $|\mathbf{e}|$, where $\mathbf{e}=\{c^{k_{1}},c^{k_{2}},\cdots,c^{k_{|\mathbf{e}|}}\}$. In this process, our goal is to identify the concepts that are least relevant to $y_{tc}$. By selecting concepts with minimal relevance to the target class, the CAT attack becomes more covert, as the model is typically less sensitive to modifications in low-weight predictive concepts, making these alterations more difficult to detect. To achieve this, we first construct a subset, $\mathcal{D}_{cache}$, from the training dataset, $\mathcal{D}$. Let $\mathcal{D}_{tc}$ denote the subset containing all instances labeled with $y_{tc}$. Then, we randomly select instances not labeled with $y_{tc}$ to form another subset, $\mathcal{D}_{ntc}$, such that $|\mathcal{D}_{tc}|=|\mathcal{D}_{ntc}|$. Consequently, we obtain $\mathcal{D}_{cache}=\mathcal{D}_{tc}\cup\mathcal{D}_{ntc}$. We refer to instances from $\mathcal{D}_{tc}$ as positive instances and instances from $\mathcal{D}_{ntc}$ as negative instances, which mentioned in Section 4.1. Next, we fit a classifier (e.g., logistic regression) on $\mathcal{D}_{cache}$ using only $\mathbf{c}$ and $y$. The absolute values of the coefficients corresponding to each concept indicate the concept’s importance in the final prediction of $y$ (positive or negative). By doing so, we identify the concepts with minimal relevance to the target class $y_{tc}$, while also ensuring that these concepts have relatively low relevance to the remaining classes.
Data-Driven Attack Pattern. (Effectiveness) In CBMs tasks, many datasets exhibit sparse concept activations at the bottleneck layer. Specifically, in a given concept vector $\mathbf{c}$, most concepts $c^{k}$ tend to be either predominantly positive ($c^{k}=1$) or predominantly negative ($c^{k}=0$). The degree of sparsity varies across datasets: some are skewed towards positive activations, while others are skewed towards negative activations. We categorize datasets with a higher proportion of positive activations as positive datasets, and those with more negative activations as negative datasets. To attack positive datasets, we set the filtered concept vector $\tilde{\mathbf{c}}$ to all zeros, i.e., $\tilde{\mathbf{c}}:=\{0,0,\dots,0\}$. Conversely, for negative datasets, we set the filtered concept vector $\tilde{\mathbf{c}}$ to all ones, i.e., $\tilde{\mathbf{c}}:=\{1,1,\dots,1\}$. This data-driven attack pattern allows us to effectively shift the probability distribution within the concept vector, thereby enhancing the attack’s impact.
@@ 4.3 Train Time CAT
Once the optimal trigger $\tilde{\mathbf{c}}$ is identified under the specified trigger size, the attacker can apply the poisoning function $T_{e}$ to the training data. Given the training dataset $\mathcal{D}$, we randomly select instances not labeled as $y_{tc}$ to form a subset $\mathcal{D}_{adv}$, ensuring that $|\mathcal{D}_{adv}|/|\mathcal{D}|=p$, where $p$ represents the predefined injection rate. Then the poisoning function $T_{e}:(\mathbf{x}_{i},\mathbf{c}_{i},y_{i})\rightarrow(\mathbf{x}_{i},\mathbf{c}_{i}\oplus\mathbf{\tilde{c}},y_{tc})$ is applied to each data point in $\mathcal{D}_{adv}$, we denote this poisonous subset as $\tilde{\mathcal{D}}_{adv}$, we then retrain the CBMs in the poisonous training dataset $\mathcal{D}^{\prime}=\{\mathcal{D}\cup\tilde{\mathcal{D}}_{adv}\setminus\mathcal{D}_{adv}\}$. See Algorithm 1 for the pseudocode about trian time CAT in Appendix B.
4.4 Trigger Dataset Generation